Calculate the simple interest and the maturity value of each of the following loans. A. Principal=$2000, Rate=7.2%, time =9 months B. Principal=$1500, rate=5.75%, time=6 months
Answer #1

simple interest=Principal*Interest Rate*Time period

A:

simple interest=(2000*7.2%*(9/12)=$108

Maturity value=simple interest+Principal

(2000+108)=$2108

B:

simple interest=(1500*5.75%*(6/12)=$43.125

Maturity value=simple interest+Principal

(1500+43.125)=$1543.125